OpenStudy (daenio):

A farmhand does 972 J of work pulling an empty hay wagon along level ground with a force of 310 N [23 below the horizontal]. A frictional force of 280 N opposes the motion. The distance the wagon travels is a. 0.39 m d. 32 m b. 1.7 m e. 1.8  102 m c. 3.4 m

OpenStudy (anonymous):

The force applied by the farmhand is F = 310N cos(theta) Then use: Work = F * d to find the distance

OpenStudy (daenio):

972J= 310Ncos31*d

OpenStudy (daenio):

I'm not sure how to do it after that.

OpenStudy (daenio):

@ten

OpenStudy (anonymous):

W = F d d = W/F Isn't your angle 23 degrees? I'm not picturing 'below the horizontal. Do you have a picture of where that angle is?

OpenStudy (daenio):

I do not.

OpenStudy (anonymous):

Do you understand: d = W/F that should give you your answer

OpenStudy (daenio):

BUT which W value?

OpenStudy (daenio):

There's 310 N [23 below the horizontal] and A frictional force of 280 N. :P Which one?

OpenStudy (anonymous):

W stands for work. It's the first thing given in the problem. 972J

OpenStudy (anonymous):

The frictional force just requires the farmer to exert more force to get the wagon to move. Only the force the farmer exerts determines how much work he does.